# CogAD: Cognitive-Hierarchy Guided End-to-End Autonomous Driving

### v2

@@ 4.2 Ablation Study
We validate the effectiveness of model design details, including the usage of BEV adapter, Mode Sharing, and Ego-Motion interaction. Mode Sharing indicates whether the ego query and motion query utilize distinct or shared trajectory mode embeddings. In the Ego-Motion interaction ablation, we remove the bidirectional interaction between ego query and motion query, retaining only the one-way interaction where ego query cross-attends motion query, which destroys the integrity of interactions between instances. As shown in Table 5, BEV adapter significantly improves the planning performance, which decreases the collision rate by nearly 90%. The removal of either Mode Sharing or bidirectional Ego-Motion interaction maintains comparable L2 error in CogAD, but elevates the collision rate from 0.06% to 0.13% (116.6% rise).
@@ 4.3 Qualitative Results
Figure 4: Qualitative results of CogAD. The 2nd column shows the Top-3 multi-mode trajectories of the highest-probability intent, with colors of red, orange, and yellow, respectively. The 3rd column displays the Top-10 multi-intent trajectories with the highest-probability mode. GT trajectory is drawn in green.
We visualize the planning results of CogAD in Figure 4. Both the surrounding camera images and the prediction results in BEV space are provided accordingly. CogAD exhibits the capability to generate hierarchically dual-level multimodal trajectories. The rightmost column displays the highest-confidence trajectories corresponding to the top-10 intents, demonstrating that intent-level multimodal planning is diverse yet valid. The second column illustrates trajectory-level multi-mode planning conditioned on the dominant intent, with visualizations of the top-3 candidate trajectories. Multi-mode planning produces multiple refined trajectory options that maintain motion consistency. By combining multi-intent diversity and multi-mode refinement, CogAD generates multiple behaviorally diverse yet physically plausible planning candidates. More visualizations appear in the supplemental material.
@@ 5 Conclusion
Limitations: Despite demonstrating excellent performance in extensive evaluations, real-world driving tests remain necessary for further validation. Additionally, the effect of model scaling requires further investigation to potentially enhance CogAD’s performance.
@@ Appendix A Datasets
We follow TOKEN [49] and manually construct a long-tail scenario validation set curated from nuScenes for comprehensive evaluation, including three scenarios: 1) executing 3-point turns; 2) resuming motion after a full stop; 3) overtaking parked cars through the oncoming lane. Details of each scenario are provided in Table 6.
@@
Closed-Loop.
Appendix B Implementation Details
Appendix C More Qualitative Results
We provide more visualization results to illustrate the effectiveness of CogAD on various driving scenarios as shown in Figure 5 and Figure 6.

### v3

@@ 4.2 Ablation Study
We validate the effectiveness of model design details, including the usage of BEV adapter, Mode Sharing, and Ego-Motion interaction. Mode Sharing indicates whether the ego query and motion query utilize distinct or shared trajectory mode embeddings. In the Ego-Motion interaction ablation, we remove the bidirectional interaction between ego query and motion query, retaining only the one-way interaction where ego query cross-attends motion query, which destroys the integrity of interactions between instances. As shown in Tab.˜5, BEV adapter significantly improves the planning performance, which decreases the collision rate by nearly 90%, demonstrating its clear advantage over using shared BEV features alone. The removal of either Mode Sharing or bidirectional Ego-Motion interaction maintains comparable L2 error in CogAD, but elevates the collision rate from 0.06% to 0.13% (116.6% rise).
Figure 5: Qualitative results of CogAD on Bench2Drive.
@@ 4.3 Qualitative Results
In Fig.˜4, we visualize the planning results of CogAD on nuScenes. Both the surrounding camera images and the prediction results in BEV space are provided accordingly. CogAD exhibits the capability to generate hierarchically dual-level multimodal trajectories. The rightmost column displays the highest-confidence trajectories corresponding to the top-10 intents, demonstrating that intent-level multimodal planning is diverse yet valid. The second column illustrates trajectory-level multi-mode planning conditioned on the dominant intent, with visualizations of the top-3 candidate trajectories. Multi-mode planning produces multiple refined trajectory options that maintain motion consistency. By combining multi-intent diversity and multi-mode refinement, CogAD generates multiple behaviorally diverse yet physically plausible planning candidates.
Our qualitative analysis demonstrates that CogAD exhibits behaviors analogous to human cognition, as evidenced by the highest-probability planned trajectory visualized in Fig.˜5 and the video recordings in the Supplementary Material. These behaviors include strategic speed reduction for yielding, proactive lane changes for obstacle avoidance, and context-aware trajectory planning.
@@ A Datasets
We follow TOKEN [49] and manually construct a long-tail scenario validation set curated from nuScenes for comprehensive evaluation, including three scenarios: 1) executing 3-point turns; 2) resuming motion after a full stop; 3) overtaking parked cars through the oncoming lane. Details of each scenario are provided in Tab.˜6.
Closed-Loop.
@@
Figure 8: Qualitative results of CogAD on nuScenes.
Figure 9: Qualitative results of CogAD on Bench2Drive.
B Implementation Details
C More Qualitative Results
We provide more visualization results to illustrate the effectiveness of CogAD on various driving scenarios as shown in Fig.˜6, Fig.˜7, Fig.˜8, and Fig.˜9.